OpenStudy (anonymous):

Please could use some help Translate the sentence into an equation. The difference of a number times 3 and 6 is equal to 5. Use the variable w for the unknown number.

OpenStudy (anonymous):

The Difference = Subtraction . Soo ... 3w ( Or In Your Case, A Number ) - 6 = 5 .

OpenStudy (anonymous):

3 Times A Number = 3w ( W = A/Any Number )

OpenStudy (anonymous):

So 3w(w=a)-6=5

OpenStudy (anonymous):

No I Was Saying w = A Number OR Any Number . Basically The Same Statement I'm Just Rewording It . But In Your Question Where It Says 3 Times A Number . As You See Later In That Question It Says Use "w" As The Unknown Number. w = ( A Number ) . Thus Making The First Part Of The Question "3w" . Do You Get What I'm Saying ?